40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2426     Entry time: Fri Sep 13 16:46:54 2019
Author: anchal 
Type: Summary 
Category: TempCtrl 
Subject: Cavity heater resistances remeasured 

Present understanding

  • CTN:2035 was the last measurement of cavity shield heater resistances. These were reported as 85.6 Ohm and 156.8 Ohm.
  • So Andrew, undestandably, configured the EPICs channel such that to put in 1 W power on the south heater, more current was driven through it in comparison to North heater.
  • This should have worked well, but over the past year atleast, I have been noticing that south heater is much more effective than the north one.

My hypothesis

  • I remeasured the resistances and they came almost same: South 88.4 Ohm and North was 159.7 Ohm.
  • It could have been the case that North side was some parasitic resistance (due to bad connection with the heater) in series, increasing its effective resistance to 159.7 Ohm.
  • This would mean that the heating part of the resistance is still almost same as South, but because we sense it larger, we send less current on this coil.
  • I assume Tara or whoever put on these heaters, they made sure the length of the wires were almost same, so it is a fairly good assumption that the heating part of resistance is same.
  • This also matches with the observation. North side is less effective in heating, because we send less current to it.

Changes made today

  • I changed the dB files for the EPICs channels, so that same 88.4 Ohm resistance is assumed for both North and South heaters.
  • From the performance so far, and looking at the second derivative of the beatnote signal, I feel the heaters are more balanced now.
  • This should help the PID as the actuation is not biased now.
  • I have made the hard actuation limits same on both ends of PID at 1.1 W.
  • We might need to retune the PID to get new PID constants.
ELOG V3.1.3-